In the laboratory, a general chemistry student measured the pH of a 0.480 M aqueous solution of quinoline, C9H7N to be 9.254.

Use the information she obtained to determine the Kb for this base.
1
Expert's answer
2017-11-23T06:24:07-0500
There is next equilibrium equation :
C9H7N + H2O ↔ C9H7NH+ + OH-
pH+pOH=14, so pOH=14-pH=14-9.254= 4.746
[OH-]= [C9H7NH+] = 〖10〗^(-4.746) =1.795∙10-5 M
K_b=([C_9 H_7 NH^+ ][OH^-])/[C_9 H_7 N] =([〖1.795∙10〗^(-5) M][〖1.795∙10〗^(-5) M])/[0.480 M] =〖6.7126∙10〗^(-10)

Answer: 6.7126∙10-10